The masses and the  radii of the   earth  and moon are M1,R1 and  M2 and R2  respectively  . their centres are  distance d apart . the minimum velocity with which aparticle  of mass m  should be proojected   froma point mdiway  between their centres  so that it escapes to   infinity is?


Grade:10

1 Answers

Souptik Naha
4 years ago
According to the conservation of energy....KE and PE must remain same....so....
(1/2)mv2 + PEearth+ PEmoon=0
(1/2)mv2 -(GM1m)/(d/2)-(GMm)/(d/2) =0
(1/2)v2=2G/d(M1+M2)
V=2 (G(M1+M2)/d)(1/2)